1. Evaluate each of the following limits: (4 pts ea) b) lim 2- a) lim VF-1 bilim 7-x
Answer #1

lim 5x - 1 - x-1 Uy Using l Hospitals Rule - M 21 X U-t is negative when xoti Therefore, 19-11= -x+7 - lim ( 17 at Xo1 ( - =